In the equation (𝑃+π‘Ž/𝑉2)(π‘‰βˆ’π‘)=π‘π‘œπ‘›π‘ π‘‘π‘Žπ‘›π‘‘, determine the units of a and b.


1
Expert's answer
2021-03-14T19:16:13-0400

Only the quantities with the same units can be added or subtracted. Thus, from the first term obtain:


"[P] = \dfrac{[a]^2}{[V]}"

where "[P] = Pa, [V] = m^3", and "[a]" is the units of "a". Thus, obtain:


"[a] = \sqrt{[P][V]} = \sqrt{Pa\cdot m^3}"

Similarly, from the second term obtain:


"[V] = [b]\\
[b] = m^3"

Answer. Units of "a" are "\sqrt{Pa\cdot m^3}", units of "b" are "m^3".